i cant understand algebra can you help me ? 9(d-1)+d=11
Answer by edjones(8007)

9(d-1)+d=11
9d-9+d=11 multiply 9 times d-1
10d-9=11
10d=20 add 9 to each side
d=2 divide each side by 10
.
Ed